In how many can 3 men and 2 women be chosen from 7 men and 5 women
Answer by Alan3354(69443)

In how many can 3 men and 2 women be chosen from 7 men and 5 women
--------
For the men, it's 7*6*5 = 210
For the women, it's 5*4 = 20
210*20 = 4200 ways